OpenStudy (anonymous):

calculate the mole fraction of benzene in solution containing 30% by mass in carbon tetra chloride

OpenStudy (jfraser):

if the solution is 30% by mass, then if you assume you have 100g of solution, how much is benzene, and how much is the CCl4?

OpenStudy (anonymous):

benzene 30 g ccl4 -70g ???

OpenStudy (jfraser):

so now, since you know grams of each, can you find moles of each?

OpenStudy (anonymous):

c6h6 -.384 mol ccl4 - .454mol

OpenStudy (jfraser):

since you know the moles of each, what is the mole fraction of benzene?

OpenStudy (anonymous):

Mass of benzene = 30 molar mass of benzene = 78 moles of benzene = 30/78 = 0.38 mass of carbon tetrachloride = 70 molar mass of carbon tetrachloride = 154 moles of carbon tetrachloride = 70/154 = 0.45 total number of moles = 0.38 + 0.45 = 0.83 mole fraction of benzene = 0.38 / 0.83 = 0.45

OpenStudy (anonymous):

http://www.meritnation.com/discuss/question/2416152

OpenStudy (anonymous):

@JOYAL did u understand???